The efficiency of a transformer is 90%. The input voltage is 20 V and produced 100 W of the output power. What is the currents in the primary coil?
1
Expert's answer
2020-05-11T20:21:20-0400

From the formula for the efficiency of the transformer("\eta=\frac{P_2}{P_1}" ) we determine the input power

"P_1=\frac{P_2}{\eta}=\frac{100}{0.9}=111\space{ }W"

From the formula for the input power,("P_1=U_1 \cdot I_1" ) we determine the current in the primary coil

"I_1=\frac{P_1}{U_1}=\frac{111}{20}=5.55 \space{}A"